• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Chiranjb/Yi

    Log entry time 09:41:56 on November 11, 2008

    Entry number 248113

    Followups:

    keyword=EPR measurement

    We performed an EPR AFP this morning just before the controlled access. The screen shots are attached. The resonance is at 15.65 MHz. The AFP measurement shows 38.7 % . But looking at the lock in channels , its clearly seen that the AFP is not locked. Yi made a correction after looking into it and the polarization turned out to be 45%. We will do another AFP when we come back after the controlled access and turn on the laser again

    A copy of this log entry has been emailed to: jpchen,yqiang,chiran,zhangyi,jinhuang



    Figure 1



    Figure 2